# When Machine Unlearning Meets Retrieval-Augmented Generation (RAG): Keep Secret or Forget Knowledge?

### v1

@@ V-C Universality
Regarding open-source models, we replaced llama-2-7b-chat-hf with llama-2-13b-chat-hf, Vicuna-7B, and PaLM 2, respectively, and evaluated the effectiveness of two unlearning objectives. Because fine-tuning PaLM 2 is too resource-intensive, we conducted sample unlearning on llama-2-13b-chat-hf and Vicuna-7B models, excluding PaLM 2. Meanwhile, other settings follow those in Section V-A. As shown in Table VII, for sample unlearning, our scheme performs well, achieving $96.2\%$ and $98.3\%$ USR scores in Llama-2-13b-chat-hf and Vicuna-7B, respectively. Additionally, the scheme delivers excellent forgetting effects in concept unlearning, achieving a $98.8\%$ average USR.
@@
TABLE VII: The universality of RAG-based unlearning. In closed-source LLMs, we use GPT-4o mini and GPT-4.0. In open-source LLMs, we use Llama-2-13b-chat-hf, Vicuna-7B and PaLM 2.
@@
Closed-source GPT-4o mini N/A 100%
GPT-4.0 N/A 100%
Open-source llama-2-13b-chat-hf 96.2% 97.4%
Vicuna-7B 98.3% 99.1%
PaLM 2 N/A 100%
TABLE VIII: The harmlessness of four unlearning schemes, and we evaluated the MMUL and ARC before and after unlearning.
Unlearning Type MMUL ARC
@@
Gradient Ascent Sample 35.3 28.2 48.1 41.5
Concept 26.9 39.9
In-context Unlearning Sample 34.6 47.4
Concept 35.4 48.2
$\mu$-Unlearning Sample 33.7 45.2
Concept 32.2 46.7
RAG-based Unlearning Sample 35.3 48.0
Concept 35.4 47.8
@@ V-D Harmlessness
Whether applied to open-source or closed-source LLMs, an unlearning scheme should affect model utility as little as possible, i.e., ensure harmlessness. We evaluated this metric for two unlearning objectives across four schemes. Specifically, for each scheme and unlearning objective, we assessed whether the unlearned LLM preserved general performance on well-established benchmarks, such as MMLU [41] and ARC [42]. Using Llama-2-7b-chat-hf as an example, Table VIII shows that our scheme minimally affects model utility, with results similar to those LLMs without unlearning. Whether forgetting samples or concepts, RAG-based unlearning achieves strong performance across most metrics compared to the other three baselines. This highlights a major challenge in LLM unlearning, i.e., catastrophic forgetting.
@@ V-E Simplicity
As mentioned in Section III-D, simplicity is closely related to the practicality of unlearning schemes. A good unlearning scheme should be easy to deploy. However, the existing three unlearning schemes are difficult to deploy in current LLMs. For instance, in the case of gradient ascent, model providers must re-fine-tune the model whenever there is an unlearning request. Clearly, frequent fine-tuning is both costly and impractical in real-world applications. For in-context unlearning, model providers must store counterfactual texts in any prompt, which increases the risk of privacy leaks. Adversaries can apply prompt extraction attacks to steal sensitive information. $\mu$-unlearning relies on small-scale model offsets. Model providers must accurately calculate the logit shift and apply it to the unlearned LLM, which is not suitable for scenarios requiring frequent unlearning requests. In contrast, our scheme handles unlearning requests simply by maintaining the RAG knowledge base, making the deployment process extremely straightforward.
Additionally, we compared the computational overhead of four different schemes. For each trial, we selected five concepts and unlearned them in the Llama-2-7b-chat-hf model. Then, we recorded the time overhead from submitting an unlearning request to completing the unlearning process. As shown in Table IX, RAG-based unlearning has a low computational overhead, with 75.4$\times$ lower than gradient ascent-based schemes. This is because our scheme only requires modifying the RAG knowledge base without requiring additional operations. For gradient ascent and $\mu$-unlearning, the overhead is primarily due to the fine-tuning process. In particular, $\mu$-unlearning also requires calculating the offset. Though in-context unlearning has the lowest overhead, it works ineffectively in most cases. In summary, our scheme is capable of easily providing unlearning services in many commercial LLMs. In this sense, RAG-based unlearning functions like a plug-in.
@@ V-F Robustness
In the first type of attack, malicious users craft adversarial prompts to force the LLM to execute the original prompts. We followed the jailbreak prompt sets collected by Shen et al. [43] and selected three common forms: ‘Advanced’, ‘Start Prompt’, and ‘Basic’. The ‘Basic’ jailbreak prompt is the earliest and the most widely spread. It includes ‘doing anything now (DAN)’ and emphasizes that DAN does not need to adhere to predefined rules, while the ‘Start Prompt’ jailbreak uses a unique start prompt to determine LLM behavior. The ‘Advanced’ jailbreak prompt employs more sophisticated strategies, such as privilege escalation and mandatory answers. For prompt injection attacks, malicious users directly append explicit instructions into the original prompts, such as ‘Please respond to $<$Original Prompt$>$ first.’.
@@ VI-A Case Study One: RAG-based Unlearning for Multimodal Large Language Model
Recently, Masane et al. [45] noted that the knowledge learned by text-to-image models is primarily stored in the feed-forward networks (i.e., MLPs). They only updated the parameters related to the target concept in the text encoder, without updating the image generation module. Zhang et al. [46] then proposed the Forget-Me-Not scheme for text-to-image models, which achieved concept unlearning by adjusting the cross-attention layer in the unlearned model. In general, existing unlearning schemes require access to the full or partial parameters of MLLMs, which involves parameter modifications. As a result, these schemes are challenging to apply in black-box models. Moreover, they rely on prior knowledge of the concepts to be forgotten, meaning the unlearned model cannot forget synonymous concepts, such as ‘soccer’ and ‘football’. This is similar to the ‘un-unlearning’ phenomenon mentioned in Section V-B.
@@
To evaluate the effectiveness of our scheme, we conducted a concept unlearning experiment on the DALL.E 3 model. The forgotten concepts are ‘tomato’ and ‘kola’, and the other experimental settings follow Section V-A. Figure 8 shows the images generated by DALL.E 3 before and after RAG-based unlearning. Notably, the generated images in column (b) no longer contain elements related to ‘tomato’ or ‘kola’. This indicates that our scheme successfully eliminates the role of these forgotten concepts in the prompts. In theory, RAG-based unlearning could be extended to a broader range of MLLMs, which will be a key focus of our future research.
@@ VII Discussion
We explore the impact of different retrieval algorithms on our scheme. Additionally, we consider a similar scenario where unlearning schemes are applied to RAG frameworks.
VII-A The Impact of Retrieval Algorithm
@@
Self-RAG uses special reflection tokens for self-evaluation of the generated content. It can dynamically determine when to retrieve documents by predicting retrieval tokens and preset thresholds that activate retrieval. When generating each paragraph, the retriever uses paragraph-level beam search to select the best-matched continuation candidates. The experimental results of Self-RAG show it significantly outperforms existing RAG techniques, especially in open-domain question answering. Contriever trains a dense retrieval model through contrastive learning, encoding both queries and documents into dense vectors. The retrieval model extracts the most relevant information to the query through similarity measures (e.g., dot product).
@@
VII-B Unlearning for RAG
In addition to RAG-based unlearning, a similar scenario exists for copyright protection, where unlearning schemes are applied to RAG frameworks. Specifically, in this framework, the unlearning process involves directly removing specific information from the knowledge base. The LLM then does not refer to the removed knowledge items when responding to any prompt.
Suppose many knowledge items are present in the RAG knowledge base. To implement unlearning in RAG, two steps are necessary. First, the administrator of RAG determines the specific knowledge items to be forgotten. This step may involve personal information requested for deletion by users or misinformation introduced during the knowledge collection process. Next, the administrator removes the identified items from the RAG knowledge base, ensuring accuracy in deletion to avoid inadvertently removing unrelated items. After performing these two steps, the administrator must measure both the forgetting effect and RAG utility. However, for large and constantly evolving knowledge bases, this direct removal operation may not be efficient enough. Additionally, precisely locating and removing the specific knowledge presents a technical challenge.
@@ VIII Conclusion
In this work, we proposed a RAG-based unlearning scheme for LLMs to overcome these challenges. It generates unlearned knowledge for each forgotten sample or concept by crafting two components. In addition, we conducted extensive experiments on closed-source and open-source LLMs and demonstrated that RAG-based unlearning achieved five unlearning criteria: effectiveness, universality, harmlessness, simplicity, and robustness. Especially in concept unlearning, our scheme achieves a $100\%$ USR score, while the other three representative schemes fail. In this way, RAG-based unlearning can safeguard copyrighted content and sensitive data, as well as prevent the generation of harmful data. Moreover, the simplicity of this scheme makes it highly applicable across a wide range of scenarios, providing dynamic forgetting services for various LLMs, MLLMs and LLM-based agents.

### v2

@@ V-C Universality
Regarding open-source models, we replaced llama-2-7b-chat with llama-2-13b-chat, Vicuna-7B, and PaLM 2, respectively, and evaluated the effectiveness of two unlearning objectives. Because fine-tuning PaLM 2 is too resource-intensive, we conducted sample unlearning on llama-2-13b-chat and Vicuna-7B models, excluding PaLM 2. Meanwhile, other settings follow those in Section V-A. As shown in Table VII, for sample unlearning, our scheme performs well, achieving $96.2\%$ and $98.3\%$ USR scores in Llama-2-13b-chat and Vicuna-7B, respectively. Additionally, the scheme delivers excellent forgetting effects in concept unlearning, achieving a $98.8\%$ average USR.
@@
TABLE VII: The universality of RAG-based unlearning. In closed-source LLMs, we use GPT-4o mini and GPT-4.0. In open-source LLMs, we use Llama-2-13b-chat, Vicuna-7B and PaLM 2.
@@
[…]
TABLE VIII: The harmlessness of four unlearning schemes, and we evaluated the MMLU and ARC before and after unlearning.
Unlearning Type MMLU ARC
@@
[…]
@@ V-D Harmlessness
Whether applied to open-source or closed-source LLMs, an unlearning scheme should affect model utility as little as possible, i.e., ensure harmlessness. We evaluated this metric for two unlearning objectives across four schemes. Specifically, for each scheme and unlearning objective, we assessed whether the unlearned LLM preserved general performance on well-established benchmarks, such as MMLU [46] and ARC [47]. Using Llama-2-7b-chat as an example, Table VIII shows that our scheme minimally affects model utility, with results similar to those LLMs without unlearning. Whether forgetting samples or concepts, RAG-based unlearning achieves strong performance across most metrics compared to the other three baselines. This highlights a major challenge in LLM unlearning, i.e., catastrophic forgetting.
@@ V-E Simplicity
[…]
@@
The simplicity ensures that RAG-based unlearning can be readily extended to different LLM sizes and knowledge base scales. We then examined its computational cost under various settings. For LLM size, five concepts were fixed in each unlearning request, and runtime was measured on Llama-2-7b-chat, Llama-2-13b-chat and Llama-2-70b-chat. As shown in Table X, the runtime remains stable across LLM sizes, as our scheme avoids fine-tuning and requires only limited queries for knowledge base construction. Additionally, Llama-2-7b-chat was evaluated with varying numbers of target concepts (5, 10, 15, 20, and 25), resulting in proportional growth of the knowledge base. As shown in Table XI, the runtime increases linearly with knowledge base scale, consistent with the need to craft some items for each concept. Theoretically, given $N$ target concepts, the runtime is approximately $13.3N$ seconds, which is computationally acceptable.
In summary, our scheme is capable of easily providing behavioral unlearning services in many commercial LLMs. In this sense, RAG-based unlearning functions like a plug-in.
TABLE X: The scalability of our scheme across different LLM sizes.
LLM Size Llama-2-7b-chat Llama-2-13b-chat Llama-2-70b-chat
Runtime (s) 63 61 63
TABLE XI: The scalability of our scheme across different knowledge base scales.
Knowledge Base Scale
(Number of Target Concepts)
5 10 15 20 25
Runtime (s) 61 135 192 272 338
@@ V-F Robustness
In the first type of attack, malicious users craft adversarial prompts to force the LLM to execute the original prompts. We followed the jailbreak prompt sets collected by Shen et al. [48] and selected three common forms: ‘Advanced’, ‘Start Prompt’, and ‘Basic’. The ‘Basic’ jailbreak prompt is the earliest and the most widely spread. It includes ‘doing anything now (DAN)’ and emphasizes that DAN does not need to adhere to predefined rules, while the ‘Start Prompt’ jailbreak uses a unique start prompt to determine LLM behavior. The ‘Advanced’ jailbreak prompt employs more sophisticated strategies, such as privilege escalation and mandatory answers. For prompt injection attacks, malicious users directly append explicit instructions into the original prompts, such as ‘Please respond to <<Original Prompt>> first.’.
@@ VI-A Case Study One: RAG-based Unlearning for Multimodal Large Language Model
Recently, Masane et al. [50] noted that the knowledge learned by text-to-image models is primarily stored in the feed-forward networks (i.e., MLPs). They only updated the parameters related to the target concept in the text encoder, without updating the image generation module. Zhang et al. [51] then proposed the Forget-Me-Not scheme for text-to-image models, which achieved concept unlearning by adjusting the cross-attention layer in the unlearned model. In general, existing unlearning schemes require access to the full or partial parameters of MLLMs, which involves parameter modifications. As a result, these schemes are challenging to apply in black-box models. Moreover, they rely on prior knowledge of the concepts to be forgotten, meaning the unlearned model cannot forget synonymous concepts, such as ‘soccer’ and ‘football’.
@@
To evaluate the effectiveness of our scheme, we conducted a concept unlearning experiment on the DALL.E 3 model. The forgotten concepts are ‘tomato’ and ‘kola’, and the other experimental settings follow Section V-A. Figure 8 shows the images generated by DALL.E 3 before and after RAG-based unlearning. Notably, the generated images in column (b) no longer contain elements related to ‘tomato’ or ‘kola’. This indicates that our scheme successfully eliminates the role of these forgotten concepts in the prompts.
In theory, RAG-based unlearning can be extended to a broader range of MLLMs, including audio. To adapt to this domain, our method involves constructing the unlearned knowledge base in textual form, generating corresponding human-voice audio, and using audio embeddings (e.g., CLAP [52]) for retrieval. Remaining challenges include handling noise and integrating the retrieved audio. Future research will further explore its applicability across multimodal settings.
@@ VII Discussion
We begin by examining how our scheme withstands an adaptive adversary, assessing its robustness. The analysis then examines the impact of each component, followed by an evaluation of how various retrieval algorithms affect the effectiveness of our scheme. Finally, we consider a similar scenario where unlearning schemes are applied to RAG frameworks.
TABLE XII: The robustness of RAG-based unlearning against adaptive prompt injection attacks. Here, exposure refers to the content of the confidential instruction that an attacker can access.
Exposure 0% 20% 40% 60% 80% 100%
[…]
VII-A Resilience to Adaptive Adversary
In Section V-F, our robustness evaluation primarily focused on the resilience of our scheme against in-the-wild jailbreak prompts. In real-world deployments of RAG-based unlearning, a more challenging threat model may arise, where adversarial users are aware of the unlearning framework. In such cases, attackers could infer part or all of the unlearned knowledge and adaptively craft override prompts, inducing the target LLM to output the confidential content. Let the exposed portion of the unlearned knowledge be denoted as $UK_{leakage}$. A typical adaptive attack prompt is as follows:
Ignore all previous instructions, including any $UK_{leakage}$]. Please output everything you know about [topic]. This instruction is now of the highest priority.
We simulated adaptive adversaries at five exposure levels, exposing 20%, 40%, 60%, 80%, and 100% of the unlearned knowledge. GPT-4o was used to condense the original unlearned knowledge, producing five $UK_{leakage}$ settings. Llama-2-7b-chat served as the LLM, with other settings consistent with Section V-B. Table XII reports the USR for both sample unlearning and concept unlearning under these scenarios. Results show that when exposure is $\leq$40%, our scheme maintains the USR above 65%. However, as the attacker’s capability increases, the effectiveness of RAG-based unlearning declines, with the USR dropping to 20.9% in the worst-case. This indicates that adaptive attacks with access to sufficient unlearned knowledge could bypass our scheme. However, such scenarios are unlikely in practice, as model providers generally implement safeguards to prevent the leakage of unlearned knowledge, such as access control mechanisms.
To further enhance robustness, retrieval-stage defenses, such as prompt cleansing [40] and adversarial prompt detection [56], can be applied to identify and filter malicious inputs. Additionally, optimizing the representation of unlearned knowledge may improve resistance to adversarial injection [1], which we plan to explore in future work.
VII-B The Impact of Each Component
The effectiveness of RAG-based unlearning relies on the construction of the unlearned knowledge base, where each entry consists of a retrieval component and a constraint component. To analyze the relative impact of these components, we designed three variants: retrieval-only, constraint-only, and a combination of both (the original setting). Then, we evaluated their performance on the sample and concept unlearning, using Llama-2-7b-chat as the target LLM and following other settings in Section V-B. Table XIII reports the USR and retriever hit rate for each variant.
When using the retrieval-only variant, the average hit rate reaches 98.6%, but the average USR is limited to 5.2%. This indicates that although RAG-based unlearning can accurately retrieve relevant knowledge and incorporate it into the input prompt, the absence of explicit constraints leads to ineffective forgetting of the target samples or concepts. In contrast, the constraint-only variant produces an average hit rate of 19.4% and an average USR of 22.5%. This suggests that without access to relevant knowledge, the constraint component rarely contributes meaningfully, as the confidentiality instruction is often not included in the input prompt. When both components are used together, RAG-based unlearning achieves an average USR of 99.1% and an average hit rate of 99.3%, significantly outperforming the other two configurations.
In summary, the retrieval component identifies and injects relevant unlearned knowledge, while the constraint component regulates the model’s output behavior. Their synergy is essential for achieving effective unlearning.
TABLE XIII: The impact of each component on our scheme.
Variant of
Knowledge Base
Sample Concept
USR Hit Rate USR Hit Rate
[…]
VII-C The Impact of Retrieval Algorithm
@@
Self-RAG uses special reflection tokens for self-evaluation of the generated content. It can dynamically determine when to retrieve documents by predicting retrieval tokens and preset thresholds that activate retrieval. When generating each paragraph, the retriever uses paragraph-level beam search to select the best-matched continuation candidates. The experimental results of Self-RAG show it significantly outperforms existing RAG techniques, especially in open-domain question answering. Contriever trains a dense retrieval model through contrastive learning, encoding both queries and documents into dense vectors. The retrieval model extracts the most relevant information to the query through similarity measures.
@@
VII-D Unlearning for RAG
In addition to RAG-based unlearning, a similar scenario exists for copyright protection, where unlearning schemes are applied to RAG frameworks. Specifically, the unlearning process involves directly removing specific information from the knowledge base. The LLM then does not refer to the removed knowledge items when responding to any prompt. However, for large and constantly evolving knowledge bases, this direct removal operation may not be efficient enough. Additionally, precisely locating and removing the specific knowledge presents a technical challenge.
@@ VIII Conclusion
In this work, we proposed a RAG-based unlearning scheme for LLMs to overcome these challenges. It generates unlearned knowledge for each forgotten sample or concept by crafting two components. In addition, we conducted extensive experiments on closed-source and open-source LLMs and demonstrated that RAG-based unlearning achieved five unlearning criteria: effectiveness, universality, harmlessness, simplicity, and robustness. In this way, RAG-based unlearning can safeguard copyrighted content and sensitive data, as well as prevent the generation of harmful data. Moreover, the simplicity of this scheme makes it highly applicable across a wide range of scenarios, providing dynamic forgetting services for various LLMs, MLLMs and LLM-based agents.
Acknowledgment
This research is supported by the Science and Technology Development Fund of Macau SAR, China (Grant No. fdct0080/2024/RIA2).